# Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS–FTIR

### v1

@@ RESULTS
Fig. 5B shows micro– and nanoplastic fragments released from the nylon tea bag into the water during standard tea preparation, which were deposited onto the NEMS chip via nanodrop casting. Smaller fragments were also observed on the membranes after sampling the water from the aged tea bags, where a different sampling method, based on aerosolization of the liquid, was used (see Fig. S4).
@@
Fig. 5C shows the NEMS–FTIR spectra of nylon tea bag leachates prepared in duplicates. A comparison with a reference ATR-FTIR spectrum of the nylon tea bag reveals nearly identical spectral features. Vertical lines highlight characteristic IR peaks corresponding to the chemical structure of PA [53, 49]. Exact peak positions can be found in the Supplementary Materials, section B. The same peaks appear in the NEMS–FTIR spectra of water samples in which individual empty nylon tea bags were immersed and subjected to accelerated aging conditions. As aging progressed, the nylon tea bags degraded, releasing an increasing amount of PA particles, reflected in the peaks’ increasing intensity as shown in the insets (Fig. 5D).
@@
The NEMS–FTIR spectra collected in both cases allowed for chemical identification of the material from which the tea bags were made and confirmation that this material was leaching into the tea water (Fig. 5C, D). The NEMS–FTIR spectra revealed characteristic IR peaks matching well with the literature [53], confirming the presence of PA, including the amide I (1641 cm-1) and amide II (1553 cm-1) bands. The amide I band primarily arises from C=O stretching vibrations in the amide group, while the amide II band results from a combination of N–H bending and C–N stretching vibrations, which are distinctive features of PA materials.
DISCUSSION
NEMS–FTIR spectroscopy presents a versatile and sensitive platform for detecting micro– and nanoplastics. This approach has the potential to address the growing challenge of plastic pollution and the increasing presence of nanoplastics in consumer products. NEMS–FTIR has demonstrated the ability to detect nanoplastic particles at picogram levels for various plastic materials and nanoparticle sizes. In addition to qualitative characterization, a quantitative analysis was performed. The mass of PS nanoparticles deposited on the NEMS chips was estimated from their measured absorbance, showing good agreement with the expected values. This demonstrates that NEMS–FTIR spectroscopy can provide sensitive detection, identification, and reliable quantification of nanoplastics.
NEMS–FTIR was also applied to detect polymeric particles released from nylon tea bags during tea preparation. The tea bag leachates were deposited directly onto the NEMS chips without prior concentration or filtration. Additionally, nylon tea bags subjected to accelerated aging in water under UV radiation and elevated temperatures were investigated. The intensity of the NEMS–FTIR peaks increased with the number of days of exposure to the artificially created environmental stress, indicating an increased release of polymer particles during aging. This suggests that prolonged exposure to environmental aging conditions significantly contributes to the release of polymeric particles into the surrounding medium. The comparison with online spectral databases, such as Open Specy [54] and Spectragryph [55], and literature, identified the released material as nylon–based PA. In contrast to previous studies [25, 26], which required steeping at least 60 tea bags in 200 mL of water, our approach relies on a single tea bag for the same volume of water, demonstrating higher sensitivity. Here, the untreated leachate was directly deposited onto NEMS chips without any pre–concentration, and PA particles were detected and identified using NEMS–FTIR.
However, like any analytical technique, NEMS–FTIR has certain limitations. While NEMS–FTIR can be used to detect very small masses of analytes, very dilute solutions do require pre–concentration. The largest droplet volume that can be drop cast on a NEMS chip is limited to approximately 500 nL. Considering a LOD of 353 pg for PS (see Table 1), the minimum detectable concentration of PS is 0.7 $\mu$g/mL for a single 500 nl drop of sample. However, in addition to usual pre–concentration techniques such as solid phase extraction, ultrafiltration, and diafiltration, the NEMS–FTIR chip allows new sample pre–concentration possibilities. Multiple droplets can be deposited consecutively on the NEMS chip with intermediate drying steps to increase the analyte mass deposited to the desired level. Sampling via nebulization can also be used to concentrate larger sample volumes onto the NEMS chip as shown for the tea bag leachates. Based on a typical nebulizer flow rate of 20 $\mu$L/min, a sampling time of 1 hour, and an aerosol capture efficiency of approximately 50 % for 100 nm particles [33], we estimate that concentrations as low as 0.6 ng/mL could be measurable. When dealing with complex matrices containing high concentrations of organic material relative to the target analyte, matrix simplification or removal may be required to remove spectral interferences or avoid overloading the NEMS chip.
One of the key strengths of NEMS–FTIR is its ability to determine the chemical identity of samples with minimal material quantities, independent of particle size, and the possibility for in–situ sample pre–concentration. Its non–destructive nature allows seamless integration with complementary techniques, such as SEM, EDX, or O–PTIR, to gain deeper insights into sample characteristics. Furthermore, it is a cost–effective method that utilizes commercially available FTIR instruments, making it highly accessible for routine monitoring and widespread use. Spectral analysis can be further supported by freely accessible online resources, providing an open–source approach to micro– and nanoplastic identification.
MATERIALS AND METHODS
@@ DISCUSSION
Serial dilutions and mixture preparation were carried out using UHPLC–MS grade water. A PIPETMAN P1000 micropipette, with matching DIAMOND D1000 TIPACK tips (Prod. No. F144059M and F171500, respectively, Gilson Incorporated, WI, USA), was used for the preparation. All dilutions were stored in hydrolytic class 3 soda–lime glass vials with snap–on lids (Prod. No. LC84.1 and LC87.1, respectively, Carl Roth GmbH, Germany).
All stock and diluted dispersions were first vortexed for 30 seconds using a Mini Vortexer (Heathrow Scientific, IL, USA) to ensure consistent distribution before proceeding with the successive dilution. PS nanoparticle dispersions were prepared at concentrations of 31.25–2000 $\mu$g/mL, while PP and PVC dispersions ranged from 31.25 to 500 $\mu$g/mL. The mixed dispersion of all three types of polymer nanoparticles was prepared in a 1:1:1 mass ratio (PS:PP:PVC). Table 2 summarizes all final concentrations used for each single and mixed sample. Blank samples, consisting of the same UHPLC–MS–grade water used for dilution procedures, underwent the same preparation process as the experimental samples.
Table 2: Characterization of nanoplastic dispersions. Concentrations of dispersions (A) and total deposited masses of nanoplastic particles (B), differentiating between single– and three–component dispersions. Empty cells indicate concentrations not used in the experiment for the corresponding dispersion.
@@
To simulate a typical tea brewing process, 200 mL of UHPLC–MS grade water was poured into each of two Erlenmeyer flasks (Duran, DWK Life Sciences GmbH, Germany), which had been pre–cleaned following the same procedure described in the "Preparation of single and mixed nanoplastic dispersions" section. Once the water reached its boiling point, the flasks were removed from the hot plate, and a single tea bag was immersed in each. Ten minutes later, the tea bags were carefully removed using clean metal tweezers, and the water was allowed to cool overnight. The flasks were covered with aluminum foil to prevent contamination during the cooling period. Blanks were prepared using the same procedure with the same UHPLC–MS grade water but without tea bags.
@@
The accelerated aging experiment was carried out in a SimTech Feutron Double Climate Chamber (73 x 77 x 102 cm, Feutron Klimasimulation GmbH, Germany) equipped with two Ultra–Vitalux lamps (300 W, Prod. No. 4008321543929, Osram, Germany) placed 30 cm apart (see Fig. S5 in the Supplementary Materials). These lamps emitted broad–spectrum UV light (UVA: 315–400 nm, UVB: 280–315 nm) with an irradiance of approximately 59.1 W/m2, simulating natural solar radiation to induce degradation.
@@
The chamber was operated according to a cyclic schedule: 8 hours of UV exposure at 60°C, followed by 4 hours in the dark at 25°C, mimicking natural day and night temperature fluctuations. This regimen, based on ISO 4892–3 [56] and ASTM G154 [57] standards, lasted 25 days, which is equivalent to approximately one year of natural aging [58]. Seven PA tea bags, each in one glass vial with 30 mL of UHPLC–MS water, were artificially aged, and individual water samples were collected after 1, 2, 5, 10, 15, 20, and 25 day(s) to monitor particle release and degradation. Simultaneously, blanks (water without tea bags) were collected with each respective tea bag sample.
@@
As shown in Fig. 1A, NEMS chips, or resonators, play a dual role as both sensing devices and sample holders, making them the core of NEMS–FTIR technology. The wafers used in this study to fabricate NEMS chips are made from silicon (Si) and SiN, with thicknesses of 380 $\mu$m and 50 nm, respectively, and an instrinsic tensile stress of approximately 50 MPa for the SiN. After the final manufacturing step, the KOH etching, a 1x1 mm2 SiN membrane is released. This membrane has a perforated area with a diameter of approximately 600 $\mu$m, consisting of 6 $\mu$m holes spaced 3 $\mu$m apart. It acts as the platform on which the sample is deposited for analysis. Two gold electrodes, each 10 $\mu$m wide, are positioned across the chip to enable electrical transduction, allowing mechanical vibrations to be converted into electrical signals, as demonstrated by Bešić et al. [37] and Luhmann et al. [30].
@@
A PIPEJET nanoDispenser (BioFluidix–Hamilton Freiburg, Germany), shown in Fig. S6A, was used to dispense the dispersions on the membrane of the NEMS chips. Polyimide capillaries with an inner diameter of 200 $\mu$m (Prod. No. PJ–20010) were used to deliver individual droplets of 20 nL. A 1 cm length of Tygon tubing (Prod. No. VM–20053–1) was attached to the capillary to facilitate sample loading. The stroke velocity was set to 90 $\mu$m/ms during droplet dispensing.
For each concentration of the dilution series droplets of 20 nL were deposited onto the center of individual NEMS chips. The total deposited mass of each sample is listed in Table 2B. Each deposition was performed in triplicate, with three NEMS chips prepared for each specific polymer mass. Nine blanks were also sampled to provide a reliable baseline for subtraction.
For the mixture of nanoplastic particles, one drop (20 nL) of pre–mixed dispersion was sampled onto a NEMS chip. Three blanks were prepared in parallel.
For the analysis of the tea bag leachates, three NEMS chips per tea bag (one experimental and one control), were prepared along with three blanks. Rather than sampling a single droplet, each NEMS chip was sequentially loaded with five drops (20 nL each) of tea bag leachates, with a two–minute delay between drops, totaling 100 nL per chip, effectively pre–concentrating the sample.
@@
A system incorporating the Portable Aerosol Generation System (PAGS, Handix Scientific Inc., CO, USA) was employed to sample particles released from the nylon tea bags subjected to accelerated weathering cycles. This setup enabled consistent aerosolization and deposition of particles onto the NEMS chips facilitated by the holes in the central region of the membrane. Sampling was conducted at regular intervals (day 1, day 2, day 5, day 10, day 15, day 20, and day 25) to monitor the temporal evolution of particle release. The NEMS chip was placed in a holder (AEROSOL FLOW ADAPTER, Invisible–Light Labs GmbH, Austria), which was then positioned in an eight–channel filter sampler (FILT, Brechtel Manufacturing Inc., CA, USA) equipped with a built–in pump. The pump was set to a flow rate of 0.5 L/min for 5 minutes for each NEMS chip, which allowed sample deposition onto the membranes through the inertial impaction mechanism [33]. Schematic of the setup is presented in Fig. S6B.
@@
A Hitachi SU8030 scanning electron microscope (SEM, Hitachi, Japan) was employed to confirm the sizes of the particles deposited on the NEMS chips and to analyze the leachates from the plastic tea bags. For imaging, the NEMS chips were mounted on an aluminum sample holder. Secondary electron (SE) detection in the upper lens mode was utilized to capture high–resolution surface details. The microscope operated at an accelerating voltage of 2 kV with an emission current of 3900 nA. NEMS chips were examined at various magnifications to assess the particle dimensions and the structural characteristics of the deposited material.
@@
During the measurement process, sampled NEMS chips (Invisible–Light Labs GmbH, Austria) were placed in the vacuum chamber of the nanomechanical IR analyzer at 10–5 mbar and the temperature of the chips was consistently regulated to 25°$\degree$°C using the integrated thermoelectric cooling (TEC) controller and Peltier element. A resolution of 4 cm-1, with a stabilization delay of 30 ms, 200 co–additions, and an aperture of 6 mm, were used for the FTIR parameters. The spectral range extended from 4000 to 400 cm-1.
Processing of NEMS–FTIR spectra
@@
Spectral calibration was carried out using the intrinsic SiN absorption peak at 835 cm-1 as a reference. This calibration was based on transmission measurements through 12 clean, perforated, $2\times 2$ mm2, 50 nm thick low–stress SiN NEMS membranes. Fig. S7 in the Supplementary Materials shows the average transmittance and corresponding absorptance of the NEMS chip’s SiN material measured using the internal detector of the FTIR spectrometer (2 mm aperture, 32 scans, and a resolution of 2 cm-1). Despite the different sizes of the membranes ($1\times 1$ mm2 and $2\times 2$ mm2) and perforation areas (600 $\mu$m and 1200 $\mu$m, respectively), the ratio of perforated to non–perforated membrane material remains consistent across both membrane types, making the calibration based on the SiN absorption peak possible.
@@
Depositing the nanoplastics on the membrane with the nanodropper created a circular sample spot (see Fig. S1 in Supplementary Materials). The nanoplastics occupied the area $\Sigma_{\text{S}}$, which is smaller than the area of the IR beam illuminating the chip, $\Sigma_{\text{IR}}$. Furthermore, the increased photothermal response, if the sample is concentrated in the membrane center instead of being evenly distributed over the entire membrane, was corrected by $\gamma$ [60]. The sample absorptance $\alpha_{\text{S}}(\tilde{\nu})$ can now be calculated from the measured NEMS–FTIR signal of the sample, $R_{\text{S}}$, as:
@@
$\gamma$ was calculated by finite element method (FEM) simulations by Kanellopulos et al. [61]. The results for a square low-stress SiN membrane with a side length of 1 mm are given in Fig. S8 in the Supplementary Materials, yielding $\gamma=1.68\pm 0.05$ for the given sample spot size. The nanoplastics spots on the membrane formed a coffee ring, as seen in (see Fig. S1 in the Supplementary Materials). Up to 90% of particles typically accumulate at the droplet perimeter in the coffee ring, as shown for nanometer-sized PS dispersions [62, 63]. FEM simulations have shown that the formation of a coffee ring has no significant effect on the NEMS–FTIR response.
@@
This framework aims to estimate the actual mass of the nanoplastic on the NEMS chip responsible for the NEMS–FTIR signal. Following the approach of Dudani and Takahama [64], it was assumed that particle scattering is negligible compared to absorption; thus, correction factors related to scattering are not required. Particles are much smaller than the wavelength of IR radiation used in this study, placing them in the Rayleigh regime, where absorption dominates IR attenuation. In this case, the absorbance $A(\tilde{\nu})$ of the deposited particles can be expressed as a function of the linear attenuation coefficient $\mu_{\text{10}}(\tilde{\nu})$ :
@@
$m=\beta\frac{R_{\text{S}}(\tilde{\nu})-R_{\text{B}}(\tilde{\nu})}{\mu_{\text{10}}(\tilde{\nu})}\frac{\rho}{\ln(10)}\frac{\Sigma_{IR}}{\gamma},$ (6)

### v2

@@ INTRODUCTION
Figure 5B shows micro- and nanoplastic fragments released from the nylon tea bag into the water during standard tea preparation, which were deposited onto the NEMS chip via nanodrop casting. Smaller fragments were also observed on the membranes after sampling the water from the aged tea bags, where a different sample deposition method, based on aerosolization of the liquid, was used (see Figure S4).
Figure 5C shows the NEMS–FTIR spectra of nylon tea bag leachates prepared in triplicates at two aliquot volumes, 100 nL and 500 nL, deposited onto NEMS chips as-is, without prior concentration steps. A comparison with the ATR–FTIR spectrum of 500 nL of the same tea bag leachate reveals that this volume was insufficient for ATR–FTIR detection, as no characteristic peaks were identified. In contrast, the reference ATR–FTIR spectrum of the nylon tea bag itself exhibits well-defined spectral features that correspond closely to those observed in the NEMS–FTIR spectra of the leachates. The high sensitivity of NEMS–FTIR not only enables the use of very small sample volumes but also significantly reduces preparation time, as no concentration is necessary. Volumes of 500 nL of leachate were deposited on the NEMS chip and dried within approximately 30 minutes using the drop casting method (see Sampling methods in Methods). Since this volume failed to produce a detectable signal with ATR–FTIR, a larger sample would be required, which in turn implies longer drying times and more time-consuming sample preparation.
The raw, uncorrected ATR-FTIR spectrum of the tea bag is shown in Figure S5A, alongside the corrected spectrum to illustrate the effects of the applied automatic ATR correction procedures. As can be observed in Figure 5C (see closeup spectrum in Figure S5B), weak diamond phonon absorption bands originating from the ATR crystal [63] appear in the 2500–1800 cm-1 region. These arise due to the refractive index contrast between the sample and diamond ($n=2.42$ [66]). Since the refractive index of nylon-based PA is around 1.50–1.53 [67], the difference ($\Delta n$) is sufficient to make the crystal phonon features visible in the spectrum [63]. Additionally, CO2 absorption peaks are visible within the same region (Figure S5B), contributing to the observed spectral features. The variations in peak position and peak shapes, as well as the presence of additional diamond absorption bands associated with ATR-FTIR spectra, CO2 peaks and potential water vapour peaks, are all absent from the NEMS–FTIR spectra.
@@
Vertical lines, as shown in Figure 5C, highlight characteristic IR peaks corresponding to the chemical structure of PA [68, 69]. The peak at 1720 cm-1 in the NEMS–FTIR spectrum likely originates from imide groups formed during thermal–oxidative degradation of nylon, induced by brewing the tea bag in 95°$\degree$°C water [70]. Exact peak positions can be found in the Supporting Information, section B. The same peaks appear in the NEMS–FTIR spectra of water samples in which individual empty nylon tea bags were immersed and subjected to accelerated aging conditions. As aging progressed, the nylon tea bags degraded further, releasing an increasing amount of PA particles, reflected in the peaks’ increasing intensity as shown in the insets (Figure 5D). This observation is in line with previous studies suggesting that prolonged exposure to environmental aging conditions contributes to the release of polymeric particles into the surrounding medium [71].
@@
The NEMS–FTIR spectra collected in both cases allowed for chemical identification of the material from which the tea bags were made and confirmation that this material was leaching into the tea water (Figure 5C, D). Characteristic IR peaks were observed, matching well with the reference ATR–FTIR spectrum of the original tea bag, as well as with the literature [68, 69], confirming the presence of PA, including the amide I (1641 cm-1) and amide II (1553 cm-1) bands. The amide I band primarily arises from C=O stretching vibrations in the amide group, while the amide II band results from a combination of N–H bending and C–N stretching vibrations, which are distinctive features of PA materials. The comparison with online spectral databases, such as Open Specy [72] and Spectragryph [73], identified the released material as nylon–based PA.
CONCLUSIONS
NEMS-FTIR has proven effective for rapidly identifying and quantifying nanoplastics in aqueous dispersions at picogram-level sensitivity. It was demonstrated that drop-cast or nebulized polymer nanoparticles could be identified from polydisperse mixtures and that the measured signals from deposited PS nanoparticles could be used to estimate the mass of PS present on the NEMS chip.
To demonstrate applicability to real-world samples, NEMS-FTIR was used to identify polymer particles released from nylon tea bags during brewing. Leachates were directly analyzed on NEMS chips without prior sample preparation or preconcentration. Compared to ATR-FTIR, NEMS-FTIR offered significantly higher sensitivity, revealing distinct nylon spectral features from as little as 100 nL of sample, whereas ATR-FTIR failed to show peaks even at five times that sample volume. Accelerated aging experiments on nylon tea bags demonstrated the monitoring of the release of polymer particles under prolonged environmental stress. Unlike ATR-FTIR, which requires complex spectral data corrections, careful crystal selection, and suffers from spectral artifacts, NEMS-FTIR simplifies data interpretation through direct absorption measurements. Measurements performed under high vacuum with a nitrogen-purged optical path minimize spectral interference from ambient CO2 and water vapor, and the SiN NEMS chip material can be used as an intrinsic internal standard. NEMS-FTIR spectra are compatible with standard transmission FTIR spectral databases, facilitating the reliable identification of unknown samples.
Taking advantage of the wide availability of FTIR instrumentation, NEMS-FTIR is ideally suited for routine monitoring and widespread adoption. Its non-destructive nature further allows seamless subsequent analysis by complementary techniques such as SEM, EDX, or O-PTIR for detailed sample characterization. The demonstrated performance and ease of integration with existing workflows highlight NEMS-FTIR as a promising tool for environmental monitoring and nanomaterials analysis in general.
METHODS
@@
Serial dilutions and mixture preparation were carried out using UHPLC–MS grade water. A PIPETMAN P1000 micropipette, with matching DIAMOND D1000 TIPACK tips (Prod. No. F144059M and F171500, respectively, Gilson Incorporated, WI, USA), was used for the preparation. All dilutions were stored in hydrolytic class 3 soda-lime glass vials with snap–on lids (Prod. No. LC84.1 and LC87.1, respectively, Carl Roth GmbH, Germany).
All stock and diluted dispersions were first vortexed for 30 seconds using a Mini Vortexer (Heathrow Scientific, IL, USA) to ensure consistent distribution before proceeding with the successive dilution. PS nanoparticle dispersions were prepared at concentrations of 31.25-2000 $\mu$g/mL, while PP and PVC dispersions ranged from 31.25 to 500 $\mu$g/mL. The mixed dispersion of all three types of polymer nanoparticles was prepared in a 1:1:1 mass ratio (PS:PP:PVC). Table 2 summarizes all final concentrations used for each single and mixed sample. Blank samples, consisting of the same UHPLC–MS grade water used for dilution procedures, underwent the same preparation process as the experimental samples.
Table 2: Characterization of nanoplastic dispersions. Concentrations of dispersions (A) and total deposited masses of nanoplastic particles (B), differentiating between single- and three-component dispersions. Empty cells indicate concentrations not used in the experiment for the corresponding dispersion.
@@
To simulate a typical tea brewing process, 200 mL of UHPLC–MS grade water was poured into an Erlenmeyer flask (Duran, DWK Life Sciences GmbH, Germany), which had been pre–cleaned following the same procedure described in the "Preparation of single and mixed nanoplastic dispersions" section. Once the water reached its boiling point, the flask was removed from the hot plate and a single tea bag was immersed. After ten minutes, the tea bag was carefully removed using clean metal tweezers, and the water was allowed to cool overnight. The flask was covered with aluminum foil to prevent contamination during the cooling period. A blank sample was prepared using the same procedure with the same UHPLC–MS grade water but without the tea bag.
@@
The accelerated aging experiment was carried out in a SimTech Feutron Double Climate Chamber (73 x 77 x 102 cm, Feutron Klimasimulation GmbH, Germany) equipped with two Ultra–Vitalux lamps (300 W, Prod. No. 4008321543929, Osram, Germany) placed 30 cm apart (see Figure S6 in the Supporting Information). These lamps emitted broad-spectrum UV light (UVA: 315-400 nm, UVB: 280-315 nm) with an irradiance of approximately 59.1 W/m2, simulating natural solar radiation to induce degradation.
@@
The chamber was operated according to a cyclic schedule: 8 hours of UV exposure at 60°C, followed by 4 hours in the dark at 25°C, mimicking natural day and night temperature fluctuations. This regimen, based on ISO 4892-3 [74] and ASTM G154 [75] standards, lasted 25 days, which is equivalent to approximately one year of natural aging [76]. Seven PA tea bags, each in one glass vial with 30 mL of UHPLC–MS water, were artificially aged, and individual water samples were collected after 1, 2, 5, 10, 15, 20, and 25 day(s) to monitor particle release and degradation. Simultaneously, blanks (water without tea bags) were collected with each respective tea bag sample.
@@
As shown in Figure 1A, NEMS chips, or resonators, play a dual role as both sensing devices and sample holders, making them the core of NEMS–FTIR technology. The wafers used in this study to fabricate NEMS chips are made from silicon (Si) and SiN, with thicknesses of 380 $\mu$m and 50 nm, respectively, and an instrinsic tensile stress of approximately 50 MPa for the SiN. After the final manufacturing step, the KOH etching, a 1x1 mm2 SiN membrane is released. This membrane has a perforated area with a diameter of approximately 600 $\mu$m, consisting of 6 $\mu$m holes spaced 3 $\mu$m apart. It acts as the platform on which the sample is deposited for analysis. Two gold electrodes, each 10 $\mu$m wide, are positioned across the chip to enable electrical transduction, allowing mechanical vibrations to be converted into electrical signals, as demonstrated by Bešić et al[42] and Luhmann et al[35].
@@
Various drop casting strategies may be applied depending on the specific requirements of sample deposition. In this work, two different drop casting approaches were employed to deposit samples onto NEMS chips, selected based on the target deposition volume.
A PIPEJET nanoDispenser (BioFluidix-Hamilton Freiburg, Germany), shown in Figure S7A, was used to dispense the dispersions on the membrane of the NEMS chips. Polyimide capillaries with an inner diameter of 200 $\mu$m (Prod. No. PJ-20010) were used to deliver individual droplets of 20 nL. A 1 cm length of Tygon tubing (Prod. No. VM-20053-1) was attached to the capillary to facilitate sample loading. Liquid samples were inserted into the tubing using a PIPETMAN P1000 micropipette with matching DIAMOND D1000 TIPACK tips (Prod. No. F144059M and F171500, respectively, Gilson Incorporated, WI, USA). The stroke velocity was set to 90 $\mu$m/ms during droplet dispensing.
For each concentration of the dilution series, droplets of 20 nL were deposited onto the center of the individual NEMS chips. The total deposited mass of each sample is listed in Table 2B. Each deposition was performed in triplicate, with three NEMS chips prepared for each specific polymer mass. Nine additional NEMS chips were sampled with blank samples to provide a reliable baseline for subtraction.
For the mixture of nanoplastic particles, a single replicate was prepared by depositing one 20 nL drop of the pre–mixed dispersion onto a NEMS chip. Blank samples were deposited in parallel onto three additional NEMS chips.
For the analysis of the tea bag leachates, three NEMS chips were sampled with leachate samples and three with blank samples using the PIPEJET nanoDispenser. Each NEMS chip was sequentially loaded with ten drops (10 nL each) of liquid sample, with a two–minute delay between drops, totaling 100 nL per chip.
Furthermore, three NEMS chips were prepared using the Drop casting Accessory (Invisible–Light Labs GmbH, Austria), a PIPETMAN P2 micropipette, and corresponding DIAMOND DL10 TIPACK tips (Prod. No. F144054M and F171200, respectively, Gilson Incorporated, WI, USA). A single 500 nL droplet of the tea bag leachate was deposited onto each chip. The chips were then left to dry inside the Drop casting Accessory, shown in Figure S7B, in an environment with a controlled humidity gradient, which, due to the presence of perforations in the chip, facilitates droplet drying within the perforated area. In parallel, three additional NEMS chips were sampled using blank samples.
@@
A system incorporating the Portable Aerosol Generation System (PAGS, Handix Scientific Inc., CO, USA) was employed to sample particles released from the nylon tea bags subjected to accelerated weathering cycles. This setup enabled consistent aerosolization and deposition of particles onto the NEMS chips facilitated by the holes in the central region of the membrane. Sampling was conducted at regular intervals (day 1, day 2, day 5, day 10, day 15, day 20, and day 25) to monitor the temporal evolution of particle release. The NEMS chip was placed in a holder (Aerosol Flow Adapter, Invisible–Light Labs GmbH, Austria), which was then positioned in an eight-channel filter sampler (FILT, Brechtel Manufacturing Inc., CA, USA) equipped with a built-in pump. The pump was set to a flow rate of 0.5 L/min for 5 minutes for each NEMS chip, which allowed sample deposition onto the membranes through the inertial impaction mechanism [38]. A schematic of the setup is presented in Figure S7C.
@@
A Hitachi SU8030 SEM (Hitachi, Japan) was employed to confirm the sizes of the particles deposited on the NEMS chips and to analyze the leachates from the plastic tea bags. For imaging, the sampled NEMS chips were mounted on an aluminum sample holder. Secondary electron (SE) detection in the upper lens mode was utilized to capture high-resolution surface details. The microscope operated at an accelerating voltage of 2 kV with an emission current of 3900 nA. NEMS chips were examined at various magnifications to assess the particle dimensions and the structural characteristics of the deposited material.
@@
During the measurement process, sampled NEMS chips were placed in the vacuum chamber of the nanoelectromechanical IR analyzer at 10-5 mbar and the temperature of the chips was consistently regulated to 25°$\degree$°C using the integrated thermoelectric cooling (TEC) controller and Peltier element. To minimize atmospheric interferences, the optical path of the EMILIE and FTIR spectrometer were purged with dry nitrogen throughout the measurements. A resolution of 4 cm-1, with a stabilization delay of 30 ms, 200 co-additions, and an aperture of 6 mm, were used for the FTIR parameters. The spectral range extended from 4000 to 400 cm-1.
ATR–FTIR spectroscopy
ATR–FTIR spectroscopy was used alongside the NEMS–FTIR to analyze the tea bag leachates as a reference IR technique. The tea bag material itself was also examined using ATR-FTIR to confirm it was the source of the detected components in the leachates. A Platinum ATR accessory with a diamond crystal (A225/Q, Bruker Optics, MA, USA) was used for all ATR measurements, performed on the same Bruker Vertex 70 spectrometer utilized for the NEMS–FTIR experiments.
To match the NEMS–FTIR experimental conditions, an aliquot of 500 nL of tea bag leachate was applied to the isopropanol–cleaned ATR diamond crystal. The droplet was dispensed using the same micropipette and tip combination as that used for sampling the NEMS chips with the same tea bag leachates (as described in the Nanoliter droplet dispensing section). Once the droplet had evaporated, the ATR measurement was performed by lowering the pressure clamp onto the remaining residue. Additionally, 500 nL of the corresponding blank sample was measured in the same manner.
A section of the tea bag material used in the brewing experiments was also analyzed. The cleaned ATR crystal was covered with the tea bag fragment, and the pressure clamp was lowered to ensure a good contact between the sample and the crystal.
Each measurement consisted of a background spectrum recorded with the empty crystal under pressure, followed by the sample spectrum (blank, leachate, or tea bag). All spectra were acquired with a resolution of 4 cm-1, over the spectral range of 4000-400 cm-1, using 32 scans, a scanner velocity of 2.5 kHz, and an aperture of 6 mm. Atmospheric compensation was enabled during all measurements to minimize the influence of the water vapor and CO2 absorption.
The ATR–FTIR spectra were processed using the "Extended ATR correction" (standard mode) and "Baseline correction" features in the OPUS software to correct for the slight distortions in peak intensity, shape, and position inherent to ATR-FTIR measurements [60, 61, 62, 63, 64]. The spectra were also smoothed using Savitzky-–Golay filtering (window length 20, polyorder 2).
Processing of NEMS-FTIR spectra
@@
Spectral calibration was carried out using the intrinsic SiN absorption peak at 835 cm-1 as a reference. This calibration was based on transmission measurements through 12 clean, perforated, $2\times 2$ mm2, 50 nm thick low–stress SiN NEMS membranes. Figure S8 in the Supporting Information shows the average transmittance and corresponding absorptance of the NEMS chip’s SiN material measured using the internal detector of the FTIR spectrometer (2 mm aperture, 32 scans, and a resolution of 2 cm-1). Despite the different sizes of the membranes ($1\times 1$ mm2 and $2\times 2$ mm2) and perforation areas (600 $\mu$m and 1200 $\mu$m, respectively), the ratio of perforated to non–perforated membrane material remains consistent across both membrane types, making the calibration based on the SiN absorption peak possible.
@@
Depositing the nanoplastics on the membrane with the nanodropper created a circular sample spot (see Figure S1 in Supporting Information). The nanoplastics occupied the area $\Sigma_{\text{S}}$, which is smaller than the area of the IR beam illuminating the chip, $\Sigma_{\text{IR}}$. Furthermore, the increased photothermal response, if the sample is concentrated in the membrane center instead of being evenly distributed over the entire membrane, was corrected by $\gamma$ [47]. The sample absorptance $\alpha_{\text{S}}(\tilde{\nu})$ can now be calculated from the measured NEMS-FTIR signal of the sample, $R_{\text{S}}$, as:
@@
$\gamma$ was calculated by finite element method (FEM) simulations by Kanellopulos et al[78]. The results for a square low-stress SiN membrane with a side length of 1 mm are given in Figure S9 in the Supporting Information, yielding $\gamma=1.68\pm 0.05$ for the given sample spot size. The nanoplastics spots on the membrane formed a coffee ring, as seen in (see Figure S1 in the Supporting Information). Up to 90% of particles typically accumulate at the droplet perimeter in the coffee ring, as shown for nanometer-sized PS dispersions [79, 80]. FEM simulations have shown that the formation of a coffee ring has no significant effect on the NEMS–FTIR response.
@@
This framework aims to estimate the actual mass of the nanoplastic on the NEMS chip responsible for the NEMS-FTIR signal. Following the approach of Dudani and Takahama [81], it was assumed that particle scattering is negligible compared to absorption; thus, correction factors related to scattering are not required. Particles are much smaller than the wavelength of IR radiation used in this study, placing them in the Rayleigh regime, where absorption dominates IR attenuation. In this case, the absorbance $A(\tilde{\nu})$ of the deposited particles can be expressed as a function of the linear attenuation coefficient $\mu_{\text{10}}(\tilde{\nu})$ :
@@
$m=\beta\frac{R_{\text{S}}(\tilde{\nu})-R_{\text{B}}(\tilde{\nu})}{\mu_{\text{10}}(\tilde{\nu})}\frac{\rho}{\ln(10)}\frac{\Sigma_{\text{IR}}}{\gamma},$ (6)